Isn't the LCM for 6,28, 48
is
two

Found 4 solutions by abcdefg, checkley71, stanbon, Nate:
Answer by abcdefg(37)

LCM= lowest common multiple
6, 12, 18, 24, 30,36, 42, 48, 54, 60, 66, 72, 78, 84 .....
28, 56, 84, .....
48, 96, 144, ..........
and keep going until you find a number thats all the same.

Answer by checkley71(8403)

IF YOU'RE LOOKING FOR A COMMON DENOMINATOR IT'S 2016
THE LOWEST COMMON FACTOR WOULD BE 2

Answer by stanbon(48551)

LCM for 6,28, 48
Least Common Multiple is the smallest number these
three will individually divide into.
To find it find the prime number factors of the three numbers, as follows:
6=2*3
28=(2^2)*7
48=(2^4)*3
Now take every different prime factor in its highest power to form the LCM,
as follows:
LCM = (2^4)*3*7= 16*21= 336
Cheers,
Stan H.

Answer by Nate(3500)

6 = 2*3
28 = 2*2*7
48 = 2*2*2*2*3
Now, choose the numbers that make six.
2*3
Now, look at the numbers that make up 28. What numbers don't we have so far?
2*3*2*7 already have one two, so do not use
Now, repeat step two for the numbers that make up 48
2*3*2*7*2*2
The product is 336